OpenStudy (swamijatin2):

Plzz help....!! Find a quadratic polynomial each whose zeroes are-3 & 2 respectively.

OpenStudy (wolf1728):

If the zeroes are -3 and 2 then the quadratic is (x+3)*(x-2)=0 x^2 +x -6=0

OpenStudy (swamijatin2):

How u done?

sam (.sam.):

With the format (x-a)(x-b)=0 where a and b are the zeros

sam (.sam.):

\[(x-a)(x-b)=0 \\ \\ (x-(-3))(x-2)=0 \\ \\ (x+3)(x-2)=0 \\ \\ ...\]

OpenStudy (swamijatin2):

OK. Then we have to just put values of zeores and just multiply it

OpenStudy (wolf1728):

Not quite. You put the values of the zeroes with OPPOSITE SIGNS. Zeroes are -3 and 2 so we enter (x+3)*(x-2)=0

OpenStudy (swamijatin2):

Ohh.... Now I understand .........can u tell me the quad. Equation whose zeroes are1&1

OpenStudy (swamijatin2):

My answer is x^2 +2x +1. But the answer is not matching with the book 's answer

sam (.sam.):

You might be doing the expansion incorrectly (x+3)(x−2), try again

OpenStudy (swamijatin2):

Find a quad. Polynomial whose zeroes are 1& 1

OpenStudy (wolf1728):

well enter the zeroes with opposite signs so we have (x-1)*(x-1)=0

OpenStudy (swamijatin2):

Then (x-1)*(x-1) X^2 -x -x +1 X^2 -2x +1

sam (.sam.):

With the expanded equation \(x^2-x-x+1\) from \((x-1)(x-1)=0\)

OpenStudy (swamijatin2):

But the answer is x^2 -x +1. In place of x^2 -2x +1

OpenStudy (wolf1728):

No the answer is x^2 -2x +1 that other answer is wrong the roots of x^2 -x +1 are imaginary (complex)

OpenStudy (swamijatin2):

Okk... Let's take another question. Zeroes are 4&1 .... Find quad . polynomial....

OpenStudy (swamijatin2):

The answer of this is x^2 -4x +1

OpenStudy (swamijatin2):

@wolf1728

OpenStudy (mww):

If a zero of a polynomial is A, then (x-A) is a factor of the polynomial. Conversely, P(A) = 0 if A is a zero of the polynomial P. (Factor theorem) So for a quadratic with zeros being A and B, the resulting factorised polynomial is (x-A)(x-B) This means you sub in your zeros A, B into the factorised form (x-A)(x-B) and expand.

OpenStudy (wolf1728):

If zeroes equal 4 and1 the quadratic is (x-4)*(x-1)=0 x^2 -5x +4 = 0